Bill and Sally go climbing together. Sally begins 30 minutes b efore Bill at 1pm. Sally averages 18 fee/min and bill averages 20 feet/min. What time will Bill catch Sally?
Answer by checkley77(12844)

18T=20(T-30)
18T=20T-600
18T-20T=-600
-2T=-600
T=-600/-2
T=300 MINUTES AFTER SALLY STARTS BILL WILL CATCH UP WITH HER.
OR 1+5=6 PM. THEY WILL MEET.
PROOF:
18*300=20(300-30)
5,400=20*330
5,400=5,400 FEET UP THE CLIMB THEY WILL MEET.